OpenStudy (anonymous):

Solving trig Equations a.) sin^2x=3 cos^2x

OpenStudy (anonymous):

Please Help!!!

OpenStudy (jtvatsim):

reposting question in better format: \[\sin^2(x) = 3\cos^2(x)\]

OpenStudy (anonymous):

I do not have a clue where to even start.

OpenStudy (jtvatsim):

well, as it stands it looks pretty bad... do you remember what sin(x)/cos(x) is equal to?

OpenStudy (anonymous):

tan x

OpenStudy (jtvatsim):

good! now, let's use that to make this question simpler. divide both sides by cos^2(x) and it starts looking better.

OpenStudy (jtvatsim):

You should get this:\[\tan^2(x) = 3\]

OpenStudy (anonymous):

So then you square root it right?

OpenStudy (jtvatsim):

yes!

OpenStudy (jtvatsim):

do not forget that you will have a positive and negative root

OpenStudy (anonymous):

ok thank you

OpenStudy (jtvatsim):

Let me know if you need any further help good luck!

OpenStudy (anonymous):

Alternatively, you can use the identity \[\sin^2x +\cos^2x=1\] Rewriting the left (or right side, appropriately), you have \[\sin^2x=3(1-\sin^2x)\\ \sin^2x=3-3\sin^2x\\ 4\sin^2x=3\\ \sin^2x=\frac{3}{4}\]

OpenStudy (anonymous):

Which gives you the difference of squares \[\sin^2x-\left(\frac{\sqrt3}{2}\right)^2=0\\ \left(\sin x+\frac{\sqrt3}{2}\right)\left(\sin x-\frac{\sqrt3}{2}\right)=0\] Easily solvable.